# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ C.1 Instance-dependent sample complexity lower bound
In this section only, we assume that the noise follows a standard Gaussian distribution : $\xi_{t}\sim\mathcal{N}\left(0,1\right)$.
@@
Denote by $q_{\theta}(.|i,j)$ the distribution of the reward when choosing action $j$ under context $i$, and by $N_{i,j}$ the number of times $(i,j)$ has been observed in the sample set. Furthermore, define $L_{T}=L_{T}(i_{1},j_{1},r_{1},\dots,i_{T},j_{T},r_{T})$ the log-likelihood ratio between two models $\mathcal{M}_{\theta}$ and $\mathcal{M}_{\theta^{\prime}}$. Then,
@@
$T\geq\frac{\operatorname{kl}\left(\delta,1-\delta\right)}{\displaystyle\inf_{\theta^{\prime}\in\operatorname{Alt}_{\varepsilon}\left(\theta\right)}\frac{1}{2}\|\theta^{\prime}-\theta\|^{2}_{\Lambda}}.$
@@
$\mathbb{E}_{\theta}[L_{T}]=\frac{T}{2}\sum_{(i,j)\in[m]\times[n]}\omega_{i,j}((\theta-\theta^{\prime})^{\top}\psi_{i,j})^{2},$
@@
$T\geq\frac{\operatorname{kl}\left(\delta,1-\delta\right)}{\frac{1}{2}(\theta^{\prime}-\theta)^{\top}\Lambda(\theta^{\prime}-\theta)}.$
Taking the infimum over every $\theta^{\prime}\in\operatorname{Alt}_{\varepsilon}\left(\theta\right)$ completes the proof.
@@
$T\geq\frac{\operatorname{kl}\left(\delta,1-\delta\right)}{2\varepsilon^{2}}\|\psi_{\pi}\|^{2}_{\Lambda^{\dagger}}.$
@@
$T\geq\frac{\operatorname{kl}\left(\delta,1-\delta\right)}{\displaystyle\inf_{\theta^{\prime}\in\operatorname{Alt}_{\varepsilon}\left(\theta\right)}\frac{1}{2}\|\theta^{\prime}-\theta\|^{2}_{\Lambda}}.$
@@
$\varepsilon\geq\|\psi_{\pi}\|_{\Lambda^{\dagger}}\frac{\sqrt{\operatorname{kl}\left(\delta,1-\delta\right)}}{\sqrt{2T}}.$
@@
$\varepsilon<\|\psi_{\pi}\|_{\Lambda^{\dagger}}\frac{\sqrt{\operatorname{kl}\left(\delta,1-\delta\right)}}{\sqrt{2T}}.$
@@
Contrary to the lower bound of Theorem 5 in (Duan et al., 2020) for linear MDPs, our lower bound does not require additional assumptions on the set of contexts.
@@ C.2 Matching finite-sample error bound
We assume in this section that $\Lambda$ is invertible, and the noise is now only assumed $1$-subgaussian. Recall that $v^{\pi}=\theta^{\top}\psi_{\pi}$ for $\psi_{\pi}=\sum_{i,j}\omega^{\pi}_{i,j}\psi_{i,j}.$ Let $\tau>0.$ Since $\psi_{\pi}$ is known, a natural estimator of $v^{\pi}$ is $\hat{v}=\hat{\theta}^{\top}\psi_{\pi}$ where
@@
is a regularized least squares estimator of $\theta$. Note that $\sum_{t=1}^{T}\psi_{i_{t},j_{t}}\psi_{i_{t},j_{t}}^{\top}+\tau I_{d}$ is invertible for any $\tau>0$ since $\sum_{t=1}^{T}\psi_{i_{t},j_{t}}\psi_{i_{t},j_{t}}^{\top}$ is a positive semi-definite matrix with nonnegative eigenvalues. Let $\widehat{\Lambda}_{\tau}:=\frac{1}{T}\left(\sum_{t=1}^{T}\psi_{i_{t},j_{t}}\psi_{i_{t},j_{t}}^{\top}+\tau I_{d}\right)$, $L=\displaystyle\sup_{i,j}\psi_{i,j}^{\top}\Lambda^{-1}\psi_{i,j}$ and $L_{\theta}=\sup_{i,j}(\psi_{i,j}^{\top}\theta)^{2}.$
@@
Assume that $T\geq 18L\log\left(12d/\delta\right)$ and $\tau\leq L\log(12d/\delta)\lambda_{\min}\left(\Lambda\right).$ With probability at least $1-\delta$,
@@
with $E_{1}:=\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{2\log(8/\delta)}$, $E_{2}:=\|\psi_{\pi}\|_{\Lambda^{-1}}\left(LL_{\theta}+(1+4\sqrt{2d})\sqrt{L}\right)\log\left(16(d+1)/\delta\right)$, $E_{3}:=\|\psi_{\pi}\|_{\Lambda^{-1}}\left(\frac{\sqrt{2}}{3}+4\sqrt{2d}+4\sqrt{L}L_{\theta}\right)L\log^{3/2}\left(16(d+1)/\delta\right)$, $E_{4}:=\|\psi_{\pi}\|_{\Lambda^{-1}}4\sqrt{2d}L^{3/2}\log^{2}\left(16(d+1)/\delta\right).$ In particular,
$\left|v^{\pi}-\hat{v}\right|\leq\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{\frac{2\log(8/\delta)}{T}}+O\left(1/T\right)$
@@
Note that this error bound matches the sampling lower bound of Theorem C.3 up to an universal constant. This result is similar to the error bound from (Duan et al., 2020), but it requires the use of variants of the classical Freedman’s inequality adapted to subgaussian random variables, which are stated in Appendix C.3. This is due to the fact that we assume in our work that the noise is subgaussian instead of bounded.
@@
Control of $L_{1}$. To control this term, which turns out to be the dominant term in the bound, we will use Proposition C.9. Let $\mathcal{F}_{t-1}=\sigma(i_{1},j_{1},r_{1},\dots,i_{t-1},j_{t-1},r_{t-1},i_{t},j_{t})$, $U_{t}=\psi_{\pi}^{\top}\Lambda^{-1}\psi_{i_{t},j_{t}},$ $e_{t}=\xi_{t}U_{t}$ and $V_{t}=\operatorname{Var}\left(e_{t}|\mathcal{F}_{t-1}\right).$ Since $\xi_{t}$ is independent of $\mathcal{F}_{t-1}$, we have
@@
where the second to last inequality follows from the $\mathcal{F}_{t-1}$-measurability of $U_{t}$ and the independence of $\xi_{t}$ from $\mathcal{F}_{t-1}$, and the last inequality follows from the fact that $\xi_{t}$ is 1-subgaussian and independent of $U_{t}$, and from $U_{t}^{2}=V_{t}.$ Hence, by Proposition C.9,
$\mathbb{P}\left(\left|L_{1}\right|\geq\varepsilon\right)\leq\mathbb{P}\left(\left|\sum_{t=1}^{T}e_{t}\right|\geq T\varepsilon\text{ and }\sum_{t=1}^{T}V_{t}\leq\sigma^{2}\right)\leq 2\exp\left(-\frac{T^{2}\varepsilon^{2}}{2\sigma^{2}}\right)+\mathbb{P}\left(\sum_{t=1}^{T}V_{t}\geq\sigma^{2}\right)$
for any $\varepsilon,\sigma>0$. we will now choose $\varepsilon$ and $\sigma$ adequately.
Note that $\sum_{t=1}^{T}V_{t}=T\psi_{\pi}^{\top}\Lambda^{-1}\widehat{\Lambda}_{0}\Lambda^{-1}\psi_{\pi}\leq T\|\psi_{\pi}\|^{2}_{\Lambda^{-1}}\|\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}\|_{{\textup{op}}}$. Indeed, since $\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}$ is a symmetric positive semi-definite matrix, the min-max theorem entails $\|\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}\|_{{\textup{op}}}=\lambda_{\max}\left(\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}\right)\geq\frac{z^{\top}\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}z}{z^{\top}z}$ for $z=\Lambda^{-1/2}\psi_{\pi}.$ Lemma B.5 from (Duan et al., 2020) consequently ensures that $\sum_{t=1}^{T}V_{t}\leq\sigma^{2}$ with probability larger than $41-\delta/$ 4 for
$\sigma^{2}=T\|\psi_{\pi}\|^{2}_{\Lambda^{-1}}(1+K_{\delta/4}).$
where $K_{\delta/4}:=\sqrt{\frac{2\log(8d/\delta)}{T}}+\frac{2\log(8d/\delta)L}{3T}.$
Then, we let $\varepsilon:=\frac{\sigma}{T}\sqrt{2\log\left(8/\delta\right)}=\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{\frac{2\log(8d/\delta)}{T}}\sqrt{1+K_{\delta/4}}$ so that
@@
Finally, $\sqrt{1+x}\leq 1+x/2$ for all $x\geq 0$ implies that $\left|L_{1}\right|\leq\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{\frac{2\log(8d/\delta)}{T}}\left(1+\frac{K_{\delta/4}}{2}\right)$ with probability larger than $21-\delta/$ 2 .
@@
where we have used $\|\Lambda^{1/2}\theta\|_{2}=\theta^{\top}\Lambda\theta=\sum_{i,j}\omega_{i,j}\left(\psi_{i,j}^{\top}\theta\right)^{2}\leq L_{\theta}$ for the last inequality.
The proof of Lemma B.6. from (Duan et al., 2020) ensures that as long as $T\geq 18L\log\left(8d/\delta\right)$ and $\tau\leq L\log(8d/\delta)\lambda_{\min}\left(\Lambda\right)$, we have $\|\Lambda^{1/2}\Delta\Lambda^{1/2}\|_{{\textup{op}}}\leq 4\sqrt{\frac{\log(8d/\delta)L}{T}}$ with probability larger than $41-\delta/$ 4. To control $\|\Lambda^{-1/2}W\|_{2}$, we first write $\Lambda^{-1/2}W=\frac{1}{T}\sum_{t=1}^{T}W_{t}$ where $W_{t}:=\Lambda^{-1/2}\xi_{t}\psi_{i_{t},j_{t}}$ defines a martingale difference sequence in $\mathbb{R}^{d}$ with respect to the filtration $(\mathcal{F}_{t})_{t\geq 0}$ that was defined earlier. We will use the concentration inequality of Proposition C.11 to perform the analysis.
Let $P_{t}:=\Lambda^{-1/2}\psi_{i_{t},j_{t}}$. Notice that $P_{t}$ is $\mathcal{F}_{t-1}$-measurable, which implies that $W_{t}=\xi_{t}P_{t}$ defines a conditionally symmetric martingale difference sequence with respect to $(\mathcal{F}_{t})$. In our case, we have $V_{1,t}=\mathbb{E}\left[W_{t}W_{t}^{\top}|\mathcal{F}_{t-1}\right]=P_{t}P_{t}^{\top}\in\mathbb{R}^{d\times d}$ and $V_{2,t}=\mathbb{E}\left[W_{t}^{\top}W_{t}|\mathcal{F}_{t-1}\right]=P_{t}^{\top}P_{t}\in\mathbb{R}$. Let us choose $Z_{1,t}:=\xi_{t}\sqrt{P_{t}P_{t}^{\top}}$ and $Z_{2,t}:=\xi_{t}\sqrt{P_{t}^{\top}P_{t}}$. By Lemma 4.3 of (Tropp, 2011b),
$\mathbb{E}\left[e^{cZ_{1,t}}|\mathcal{F}_{t-1}\right]=\mathbb{E}\left[e^{c\xi_{t}\sqrt{P_{t}P_{t}^{\top}}}|P_{t}\right]\preccurlyeq e^{c^{2}V_{1,t}/2}$
@@
$\mathbb{E}\left[e^{cZ_{2,t}}|\mathcal{F}_{t-1}\right]=\mathbb{E}\left[e^{c\xi_{t}\sqrt{P_{t}^{\top}P_{t}}}|P_{t}\right]\preccurlyeq e^{c^{2}V_{2,t}/2}.$
@@
$\max\left(\left\|\sum_{t=1}^{T}V_{1,t}\right\|_{{\textup{op}}},\left\|\sum_{t=1}^{T}V_{2,t}\right\|_{{\textup{op}}}\right)=\sum_{t=1}^{T}P_{t}^{\top}P_{t}.$
@@
$\mathbb{P}\left(\|\Lambda^{-1/2}W\|_{2}\geq\varepsilon\right)$ $\leq\mathbb{P}\left(\left\|\sum_{t=1}^{T}W_{t}\right\|_{2}\geq T\varepsilon\text{ and }\sum_{t=1}^{T}P_{t}^{\top}P_{t}\leq\sigma^{2}\right)+\mathbb{P}\left(\sum_{t=1}^{T}P_{t}^{\top}P_{t}\geq\sigma^{2}\right)$
$\leq 2(d+1)\exp\left(-\frac{T^{2}\varepsilon^{2}}{2\sigma^{2}}\right)+\mathbb{P}\left(\sum_{t=1}^{T}P_{t}^{\top}P_{t}\geq\sigma^{2}\right)$
for any $\varepsilon,\sigma>0$. We can control $\sum_{t=1}^{T}P_{t}^{\top}P_{t}$ by adapting the proof of Lemma B.9 from (Duan et al., 2020). We have
@@
By Lemma B.5. of (Duan et al., 2020), $\sum_{t=1}^{T}P_{t}^{\top}P_{t}\leq\sigma^{2}$ with probability larger than $81-\delta/$ 8 for
$\sigma^{2}=Td(1+K_{\delta/8}).$
Then, we let $\varepsilon:=\frac{\sigma}{T}\sqrt{2\log\left(16(d+1)/\delta\right)}$ so that $\mathbb{P}\left(\|\Lambda^{-1/2}W\|_{2}\geq\varepsilon\right)\leq\delta/8+\delta/8=\delta/4.$ In particular, $\|\Lambda^{-1/2}W\|_{2}\geq\sqrt{\frac{2d\log\left(16(d+1)/\delta\right)}{T}}\left(1+\frac{K_{\delta/8}}{2}\right)$ with probability smaller than $\delta/4$. Combining this result with the bound on $\|\Lambda^{1/2}\Delta\Lambda^{1/2}\|_{{\textup{op}}}$ yields that with probability larger than $1-\delta/2,$
@@
$\leq\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}}{T}\left(4\sqrt{2d\log\left(16(d+1)/\delta\right)\log\left(8d/\delta\right)L}\left(1+\frac{K_{\delta/8}}{2}\right)+\frac{\tau L_{\theta}}{\lambda_{\min}\left(\Lambda\right)}\left(1+4\sqrt{\frac{\log(8d/\delta)L}{T}}\right)\right)$
$\leq\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}}{T}\left(4\sqrt{2d\log\left(16(d+1)/\delta\right)\log\left(8d/\delta\right)L}\left(1+\frac{K_{\delta/8}}{2}\right)+LL_{\theta}\log(8d/\delta)\left(1+4\sqrt{\frac{\log(8d/\delta)L}{T}}\right)\right),$
where we used $\tau\leq L\log(8d/\delta)\lambda_{\min}\left(\Lambda\right)$ for the last inequality. In the end,

### v2

@@ C.1 Instance-dependent sample complexity lower bound
In this section only, we assume that the noise follows a centered Gaussian distribution: $\xi_{t}\sim\mathcal{N}\left(0,\sigma^{2}\right)$.
@@
Denote by $q_{\theta}(.|i,j)$ the distribution of the reward when choosing action $j$ under context $i$, and by $N_{i,j}$ the number of times $(i,j)$ has been observed in the sample set. Let $L_{T}=L_{T}(i_{1},j_{1},r_{1},\dots,i_{T},j_{T},r_{T})$ be the log-likelihood ratio between two models $\mathcal{M}_{\theta}$ and $\mathcal{M}_{\theta^{\prime}}$. Then,
@@
$T\geq\frac{\operatorname{kl}\left(\delta,1-\delta\right)}{\displaystyle\inf_{\theta^{\prime}\in\operatorname{Alt}_{\varepsilon}\left(\theta\right)}\frac{1}{2\sigma^{2}}\|\theta^{\prime}-\theta\|^{2}_{\Lambda}}.$
@@
$\mathbb{E}_{\theta}[L_{T}]=\frac{T}{2\sigma^{2}}\sum_{(i,j)\in[m]\times[n]}\omega_{i,j}((\theta-\theta^{\prime})^{\top}\psi_{i,j})^{2},$
@@
$T\geq\frac{\operatorname{kl}\left(\delta,1-\delta\right)}{\frac{1}{2\sigma^{2}}(\theta^{\prime}-\theta)^{\top}\Lambda(\theta^{\prime}-\theta)},$
and taking the infimum over every $\theta^{\prime}\in\operatorname{Alt}_{\varepsilon}\left(\theta\right)$ completes the proof.
@@
$T\geq\frac{\sigma^{2}\operatorname{kl}\left(\delta,1-\delta\right)}{2\varepsilon^{2}}\|\psi_{\pi}\|^{2}_{\Lambda^{\dagger}}.$
@@
$T\geq\frac{\sigma^{2}\operatorname{kl}\left(\delta,1-\delta\right)}{\displaystyle\inf_{\theta^{\prime}\in\operatorname{Alt}_{\varepsilon}\left(\theta\right)}\frac{1}{2}\|\theta^{\prime}-\theta\|^{2}_{\Lambda}}.$
@@
$\varepsilon\geq\|\psi_{\pi}\|_{\Lambda^{\dagger}}\frac{\sigma\sqrt{\operatorname{kl}\left(\delta,1-\delta\right)}}{\sqrt{2T}}.$
@@
$\varepsilon<\|\psi_{\pi}\|_{\Lambda^{\dagger}}\frac{\sigma\sqrt{\operatorname{kl}\left(\delta,1-\delta\right)}}{\sqrt{2T}}.$
@@
Contrary to the lower bound of Theorem 5 in (Duan et al., 2020) for linear Markov decision processes, our lower bound does not require additional assumptions on the set of contexts.
@@ C.2 Matching finite-sample error bound
We assume in this section that $\Lambda$ is invertible, and the noise is now only assumed $\sigma$-subgaussian. Recall that $v^{\pi}=\theta^{\top}\psi_{\pi}$ for $\psi_{\pi}=\sum_{i,j}\omega^{\pi}_{i,j}\psi_{i,j}.$ Let $\tau>0.$ Since $\psi_{\pi}$ is known, a natural estimator of $v^{\pi}$ is $\hat{v}=\hat{\theta}^{\top}\psi_{\pi}$ where
@@
is a regularized least squares estimator of $\theta$. Note that $\sum_{t=1}^{T}\psi_{i_{t},j_{t}}\psi_{i_{t},j_{t}}^{\top}+\tau I_{d}$ is invertible for any $\tau>0$ since $\sum_{t=1}^{T}\psi_{i_{t},j_{t}}\psi_{i_{t},j_{t}}^{\top}$ is a positive semi-definite matrix with nonnegative eigenvalues. Let $\widehat{\Lambda}_{\tau}:=\frac{1}{T}\left(\sum_{t=1}^{T}\psi_{i_{t},j_{t}}\psi_{i_{t},j_{t}}^{\top}+\tau I_{d}\right)$, $L_{\psi}:=\max_{i,j}\psi_{i,j}^{\top}\Lambda^{-1}\psi_{i,j}$ and $L_{\theta}:=\max_{i,j}|\psi_{i,j}^{\top}\theta|.$
@@
Assume that $T\geq 18L_{\psi}\log\left(8d/\delta\right)$ and $\tau\leq L_{\psi}\log(8d/\delta)\lambda_{\min}\left(\Lambda\right).$ With probability at least $1-\delta$,
@@
with $E_{1}:=\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{2\sigma^{2}\log(8/\delta)}$, $E_{2}:=\|\psi_{\pi}\|_{\Lambda^{-1}}\left(L_{\psi}L_{\theta}+\sigma(1+4\sqrt{2d})\sqrt{L_{\psi}}\right)\log\left(16(d+1)/\delta\right)$, $E_{3}:=\|\psi_{\pi}\|_{\Lambda^{-1}}\left(\frac{\sigma\sqrt{2}}{3}+4\sigma\sqrt{d}+4\sqrt{L_{\psi}}L_{\theta}\right)L_{\psi}\log^{3/2}\left(16(d+1)/\delta\right)$, $E_{4}:=\|\psi_{\pi}\|_{\Lambda^{-1}}\frac{4\sigma\sqrt{2}}{3}\sqrt{d}L_{\psi}^{3/2}\log^{2}\left(16(d+1)/\delta\right).$ In particular,
$\left|v^{\pi}-\hat{v}\right|\leq\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{\frac{2\sigma^{2}\log(8/\delta)}{T}}+O\left(1/T\right)$
@@
This error bound matches the sampling lower bound of Theorem C.3 up to an universal constant. This result is similar to the error bound from (Duan et al., 2020), but it requires the use of variants of the classical Freedman’s inequality adapted to subgaussian random variables, which are stated in Appendix C.3. This is due to the fact that we assume in our work that the noise is subgaussian instead of bounded.
@@
Control of $L_{1}$. To control this term, which turns out to be the dominant term in the bound, we will use Proposition C.9. Let $\mathcal{F}_{t-1}=\sigma(i_{1},j_{1},r_{1},\dots,i_{t-1},j_{t-1},r_{t-1},i_{t},j_{t})$, $U_{t}=\psi_{\pi}^{\top}\Lambda^{-1}\psi_{i_{t},j_{t}},$ $e_{t}=\xi_{t}U_{t}$ and $V_{t}=\operatorname{Var}\left(e_{t}|\mathcal{F}_{t-1}\right).$ We first note that for any choice of $\gamma>0,$
$\mathbb{P}\left(\left|L_{1}\right|\geq\varepsilon\right)\leq\mathbb{P}\left(\left|\sum_{t=1}^{T}e_{t}\right|\geq T\varepsilon\text{ and }\sum_{t=1}^{T}V_{t}\leq\gamma^{2}\right)+\mathbb{P}\left(\sum_{t=1}^{T}V_{t}\geq\gamma^{2}\right).$
Since $\xi_{t}$ is independent of $\mathcal{F}_{t-1}$, we have
@@
where the second to last inequality follows from the $\mathcal{F}_{t-1}$-measurability of $U_{t}$ and the independence of $\xi_{t}$ from $\mathcal{F}_{t-1}$, and the last inequality follows from the fact that $\xi_{t}$ is $\sigma$-subgaussian and independent of $U_{t}$, and from $U_{t}^{2}=V_{t}.$ Hence, by Proposition C.9,
$\mathbb{P}\left(\left|\sum_{t=1}^{T}e_{t}\right|\geq T\varepsilon\text{ and }\sum_{t=1}^{T}V_{t}\leq\gamma^{2}\right)\leq 2\exp\left(-\frac{T^{2}\varepsilon^{2}}{2\gamma^{2}}\right)$
for any $\varepsilon,\gamma>0$. Furthermore, $\sum_{t=1}^{T}V_{t}=\sigma^{2}T\psi_{\pi}^{\top}\Lambda^{-1}\widehat{\Lambda}_{0}\Lambda^{-1}\psi_{\pi}\leq\sigma^{2}T\|\psi_{\pi}\|^{2}_{\Lambda^{-1}}\|\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}\|_{{\textup{op}}}$. Indeed, since $\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}$ is a symmetric positive semi-definite matrix, the min-max theorem entails $\|\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}\|_{{\textup{op}}}=\lambda_{\max}\left(\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}\right)\geq\frac{z^{\top}\Lambda^{-1/2}\widehat{\Lambda}_{0}\Lambda^{-1/2}z}{z^{\top}z}$ for $z=\Lambda^{-1/2}\psi_{\pi}.$ Lemma B.5 from (Duan et al., 2020) consequently ensures that $\sum_{t=1}^{T}V_{t}\leq\gamma^{2}$ with probability larger than $41-\delta/$ 4 when $\gamma$ is chosen such that
$\gamma^{2}=\sigma^{2}T\|\psi_{\pi}\|^{2}_{\Lambda^{-1}}(1+K_{\delta/4})$
where $K_{\delta/4}:=\sqrt{\frac{2\log(8d/\delta)L_{\psi}}{T}}+\frac{2\log(8d/\delta)L_{\psi}}{3T}.$
Finally, we choose $\varepsilon$ adequately with $\varepsilon:=\frac{\gamma}{T}\sqrt{2\log\left(8/\delta\right)}=\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{\frac{2\sigma^{2}\log(8d/\delta)}{T}}\sqrt{1+K_{\delta/4}}$ so that
@@
Finally, $\sqrt{1+x}\leq 1+x/2$ for all $x\geq 0$ implies that $\left|L_{1}\right|\leq\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{\frac{2\sigma^{2}\log(8d/\delta)}{T}}\left(1+\frac{K_{\delta/4}}{2}\right)$ with probability larger than $21-\delta/$ 2 .
@@
where we have used $\|\Lambda^{1/2}\theta\|^{2}_{2}=\theta^{\top}\Lambda\theta=\sum_{i,j}\omega_{i,j}\left(\psi_{i,j}^{\top}\theta\right)^{2}\leq L_{\theta}^{2}$ for the last inequality.
The proof of Lemma B.6. from (Duan et al., 2020) ensures that as long as $T\geq 18L_{\psi}\log\left(8d/\delta\right)$ and $\tau\leq L_{\psi}\log(8d/\delta)\lambda_{\min}\left(\Lambda\right)$, we have $\|\Lambda^{1/2}\Delta\Lambda^{1/2}\|_{{\textup{op}}}\leq 4\sqrt{\frac{\log(8d/\delta)L_{\psi}}{T}}$ with probability larger than $41-\delta/$ 4. To control $\|\Lambda^{-1/2}W\|_{2}$, we first write $\Lambda^{-1/2}W=\frac{1}{T}\sum_{t=1}^{T}W_{t}$ where $W_{t}:=\Lambda^{-1/2}\xi_{t}\psi_{i_{t},j_{t}}$ defines a martingale difference sequence in $\mathbb{R}^{d}$ with respect to the filtration $(\mathcal{F}_{t})_{t\geq 0}$ that was defined earlier. We will use the concentration inequality of Proposition C.11 to perform the analysis.
Let $P_{t}:=\Lambda^{-1/2}\psi_{i_{t},j_{t}}$. Note that $P_{t}$ is $\mathcal{F}_{t-1}$-measurable, which implies that $W_{t}=\xi_{t}P_{t}$ defines a conditionally symmetric martingale difference sequence with respect to $(\mathcal{F}_{t})$. In our case, we have $V_{1,t}=\mathbb{E}\left[W_{t}W_{t}^{\top}|\mathcal{F}_{t-1}\right]=\sigma^{2}P_{t}P_{t}^{\top}\in\mathbb{R}^{d\times d}$ and $V_{2,t}=\mathbb{E}\left[W_{t}^{\top}W_{t}|\mathcal{F}_{t-1}\right]=\sigma^{2}P_{t}^{\top}P_{t}\in\mathbb{R}$. Let us choose $Z_{1,t}:=\xi_{t}\sqrt{P_{t}P_{t}^{\top}}$ and $Z_{2,t}:=\xi_{t}\sqrt{P_{t}^{\top}P_{t}}$. By Lemma 4.3 of (Tropp, 2011b),
$\mathbb{E}\left[e^{cZ_{1,t}}|\mathcal{F}_{t-1}\right]=\mathbb{E}\left[e^{c\xi_{t}\sqrt{P_{t}P_{t}^{\top}}}|P_{t}\right]\preceq e^{c^{2}V_{1,t}/2}$
@@
$\mathbb{E}\left[e^{cZ_{2,t}}|\mathcal{F}_{t-1}\right]=\mathbb{E}\left[e^{c\xi_{t}\sqrt{P_{t}^{\top}P_{t}}}|P_{t}\right]\preceq e^{c^{2}V_{2,t}/2}.$
@@
$\max\left(\left\|\sum_{t=1}^{T}V_{1,t}\right\|_{{\textup{op}}},\left\|\sum_{t=1}^{T}V_{2,t}\right\|_{{\textup{op}}}\right)=\sigma^{2}\sum_{t=1}^{T}P_{t}^{\top}P_{t}.$
@@
$\mathbb{P}\left(\|\Lambda^{-1/2}W\|_{2}\geq\varepsilon\right)$ $\leq\mathbb{P}\left(\left\|\sum_{t=1}^{T}W_{t}\right\|_{2}\geq T\varepsilon\text{ and }\sum_{t=1}^{T}P_{t}^{\top}P_{t}\leq\gamma^{2}\right)+\mathbb{P}\left(\sigma^{2}\sum_{t=1}^{T}P_{t}^{\top}P_{t}\geq\gamma^{2}\right)$
$\leq 2(d+1)\exp\left(-\frac{T^{2}\varepsilon^{2}}{2\gamma^{2}}\right)+\mathbb{P}\left(\sigma^{2}\sum_{t=1}^{T}P_{t}^{\top}P_{t}\geq\gamma^{2}\right)$
for any $\varepsilon,\gamma>0$. We can control $\sum_{t=1}^{T}P_{t}^{\top}P_{t}$ by adapting the proof of Lemma B.9 from (Duan et al., 2020). We have
@@
By Lemma B.5. of (Duan et al., 2020), $\sigma^{2}\sum_{t=1}^{T}P_{t}^{\top}P_{t}\leq\gamma^{2}$ with probability larger than $81-\delta/$ 8 when $\gamma$ is such that
$\gamma^{2}=\sigma^{2}Td(1+K_{\delta/8}).$
Then, we let $\varepsilon:=\frac{\gamma}{T}\sqrt{2\log\left(16(d+1)/\delta\right)}$ so that $\mathbb{P}\left(\|\Lambda^{-1/2}W\|_{2}\geq\varepsilon\right)\leq\delta/8+\delta/8=\delta/4.$ In particular, $\|\Lambda^{-1/2}W\|_{2}\geq\sqrt{\frac{2d\sigma^{2}\log\left(16(d+1)/\delta\right)}{T}}\left(1+\frac{K_{\delta/8}}{2}\right)$ with probability smaller than $\delta/4$. To ease the notation, let $C_{1}:=\log(8d/\delta)$ and $C_{2}:=\log(16(d+1)/\delta)$. Combining the previous result with the bound on $\|\Lambda^{1/2}\Delta\Lambda^{1/2}\|_{{\textup{op}}}$ yields that with probability larger than $1-\delta/2,$
@@
$\leq\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}}{T}\left(4\sqrt{2d\sigma^{2}C_{2}C_{1}L_{\psi}}\left(1+\frac{K_{\delta/8}}{2}\right)+\frac{\tau L_{\theta}}{\lambda_{\min}\left(\Lambda\right)}\left(1+4\sqrt{\frac{C_{1}L_{\psi}}{T}}\right)\right)$
$\leq\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}}{T}\left(4\sqrt{2d\sigma^{2}C_{2}C_{1}L_{\psi}}\left(1+\frac{K_{\delta/8}}{2}\right)+L_{\psi}L_{\theta}C_{1}\left(1+4\sqrt{\frac{C_{1}L_{\psi}}{T}}\right)\right),$
where we used $\tau\leq L_{\psi}C_{1}\lambda_{\min}\left(\Lambda\right)$ for the last inequality. In the end,